OpenStudy (anonymous):

solve the eqution 3x-4y=16 for x A. x=3(16+4y) B. X= 4y/3 +16/3 C. x=4y+16 D. x= 16/3-4y/3

OpenStudy (anonymous):

@hartnn can you help i think its D but im not sure

hartnn (hartnn):

3x-4y = 16 add 4y on both sides, what do u get?

hartnn (hartnn):

any problems?

OpenStudy (anonymous):

im trying to do it im just not so good with matt so i would put it like 3x-4y=16 +4y +4y ???

OpenStudy (anonymous):

then solve ?

hartnn (hartnn):

yes go ahead 3x = 4y + 16 now divide both sides by 3

OpenStudy (anonymous):

so then it would be 3x = 4y + 16 / / 3x 3x ?

OpenStudy (anonymous):

do i plus the 4y and 16 then divide by 3 ?

OpenStudy (anonymous):

sorry im really bad at math so sorry if i sound stuppid

hartnn (hartnn):

keep 4y + 16 as it is. they are not like terms, they cannot be combined and no problem, you don't sound stupid, valid questions

OpenStudy (anonymous):

okay so then would i do 3x divide by 3 and 16 divided by 3 then

hartnn (hartnn):

you mean 3x/3 , 4y /3 and 16/ 3 , right ?

OpenStudy (anonymous):

ya

hartnn (hartnn):

correct. and 3x/3 = x so we have x = 4y/3 + 16/3 makes sense?

OpenStudy (anonymous):

yes

hartnn (hartnn):

any doubts?

OpenStudy (anonymous):

thank you for explaining i mean at first i was very confused but since you helped me through the steps it was more clear to me

hartnn (hartnn):

i am glad i could help! welcome ^_^